# Learning Object Placement Programs for Indoor Scene Synthesis with Iterative Self Training

### v1

@@ Appendix A Further Details on Constraints
Figure 11. Constraint Examples: Shown are examples of constraints, their input scene and object, and their executed masks. Objects are colored by how they appear in the scene visualization. The original scene shows where the query object was originally placed. There are four masks for each constraint. Each mask represents a possible orientation of the query object. For example the align constraint contains placement options for only one orientation (the orientation of the reference wall object). The other constraints contain placement options for all possible orientations, but those locations are constrained based on the input arguments.
[…]
TODO
@@
We assume the input objects to each constraints have the following assumed properties. Objects in each category are aligned in a canonical coordinate frame. They must also come labeled with whether they are meant for holding humans (i.e. beds and chairs). Objects and their original scenes must also come with come with semantically meaningful sizes, scales, and distances since geometric heuristics described for each constraint are based on physically meaningful quantities such as the average reaching distance. Objects and scenes which satisfy this criteria come from preprocessing the 3DFRONT (Fu et al., 2021) dataset.
Shown in Figure 11 are examples of executed constraints.
Appendix B Edge Attention mechanism
@@
Figure 12. In this example, we want to place a nightstand in a room that already contains a bed and a nightstand on the left side of it. The program which specifies this placement should always predict placement on the right side of the bed. Without edge attention the program will incorrectly place the nightstand on the left side 50% of the time. With edge attention attention our model correctly attends to the spatial relationships between objects in the scene and predicts placement on the right side every time.
@@
Consider the setting in Figure 12. The partial scene contains a single bed and nightstand to the right of it. The query object is also a nightstand. Given this setting our generative model should predict a program that places the nightstand on the left of the bed 100 % of the time. Instead, the logits corresponding to which side of the bed the program will place the nightstand is split 50-50 left and right. This is likely due to nightstands appearing on either side of the bed with equal frequency. We find this is empirically true on both a model trained on real scene data and a toy setting containing just beds and nightstands. More details of this toy setting are shown in the appendix.
@@
(2) $Z=\text{Softmax}(\frac{W_{q}X(W_{k}X)^{T}}{\sqrt{d_{k}}})W_{v}X$
@@
For each object, denoted by its index $i$, we compute key and value vectors $K’$ ’ and $V’$ ’ with respect to it. The attention weights are computed as $QK^{T}+QK’^{T}$, where the $QK’^{T}$ term acts as a correction weight to the original $QK^{T}$ attention weights. The weighted sum of $V’$ ’ using these attention weights is added to the normal output of attention for the object. Given edge values $E_{i}\in\mathbb{R}^{txd}$, our edge attention mechanism adds inter object information to each output of the original attention mechanism $Z_{i}$.
(3) $Z_{i}=Z_{i}+\text{Softmax}(\frac{W_{q}X(W_{k}X)^{T}+W_{q}X(W_{ek}E_{i})^{T}}{\sqrt{d_{k}}})W_{ev}E_{i}$
Appendix C Initial Program Extraction Implementation
Here, we describe in greater detail how programs are initially extracted from scene data. As a pre-processing step, scenes with major inter-object bounding box collisions are removed from the dataset. These scenes can produce errors in the scene extraction process.
Appendix D Scene Classifier Evaluation
We wrote a simple procedural grammar for bedrooms that can generate a dataset of scenes that comes with the ground truth location distributions used to place objects. We evaluate our scene classifier in this toy setting. Every object in a scene comes with a mask representing its ”ground truth” location distribution. We use the initial program extraction to generate a program which places this object. We add this ”ground truth” mask as a positive example to the evaluation set and negative examples are generated by randomly removing constraints from the extracted program.
We found that the scene classifier can overfit to particular locations seen during training, so we employ various techniques to reduce overfitting. We employ focal loss (Lin et al., 2017) with different $\lambda$ parameters, weight examples based on the level of perturbation of the object, and employ random translation during training.
Table 2 shows the performance of different experimental choices of the scene classifier with a acceptance threshold of $0.6$. The metrics p-cons and n-cons measures for a positive and negative example, how consistently the classifier correctly labels a mask. We seek to minimize the false positive rate which is in this settings corresponds to incorrectly predicting an object as in distribution.
Table 2. We measure the quantitative performance of different design choices for the scene classifier. We wrote a procedural grammar that can generate a dataset of scenes that comes with ”ground truth” location distributions, and use those masks to create an evaluation set. We seek to minimize the false positive rate which corresponds to incorrectly predicting an object as in distribution.
method accuracy tn fp fn tp p-cons n-cons
resnet34 0.9386 0.5400 0.0131 0.0483 0.3986 0.9440 0.9160
resnet34-focal-loss-2 0.9724 0.5359 0.0172 0.0103 0.4366 0.9920 0.8960
resnet34-focal-loss-5 0.9669 0.5303 0.0228 0.0103 0.4366 0.9960 0.8760
resnet34-weighted 0.9538 0.5124 0.0407 0.0055 0.4414 0.9960 0.7720
resnet34-weighted-focal-loss-2 0.8890 0.5372 0.0159 0.0952 0.3517 0.8720 0.9080
transformer 0.9324 0.5097 0.0434 0.0241 0.4228 0.9480 0.7520
transformer-focal-loss-2 0.9538 0.5131 0.0400 0.0062 0.4407 0.9840 0.7760
transformer-focal-loss-5 0.9441 0.5221 0.0310 0.0248 0.4221 0.9680 0.8240
transformer-no-pe 0.9393 0.5290 0.0241 0.0366 0.4103 0.9760 0.8760
transformer-shift-scene-around 0.9456 0.5120 0.0435 0.0109 0.4336 0.9840 0.7560
transformer-weighted 0.9469 0.5117 0.0414 0.0117 0.4352 0.9920 0.7640
transformer-weighted-focal-loss-2 0.9634 0.5186 0.0345 0.0021 0.4448 1.0000 0.8120
Appendix E Scene annotation software
Figure 13. A screen shot of our annotation software. On the left shows a partial scene and the object hovers on the mouse pointer. Users can draw rectangles to visualize possible placements, and then confirm them once confident in their drawing.
We built a browser based scene annotation software to facilitate the annotation of partial scenes, objects, and where they could go. This was built as a web app in react. The software allows users to draw rectangles denoting possible centroid locations of the object for a given orientation. Users can visualize what these proposed placements look like in the scene before confirming these placements. We recruited 15 university students and young working professionals to participate in the annotation. No time limit was enforced, but users on average spent on 18 seconds on each partial scene and object, or 30 minutes in total for 100 scenes. A screenshot of this software are shown in Figure 13
Appendix F Baselines Details
The retraining of both baseline methods (Paschalidou et al., 2021; Ritchie et al., 2018) on 3D-FRONT (Fu et al., 2021) differ slightly from their original training setting. Both baselines are trained on the same scenes our method trains on and both without object ordering.
ATISS’s (Paschalidou et al., 2021) data preprocessing parses living rooms and dining rooms with a maximum side length of 13.2 meters, and bedrooms and libraries with a maximum side length of 6.2. Our data processing only parses scenes with maximum side of 6.2 for all room types. During our data preprocessing, scenes with significant inter-object penetration are removed from the dataset as they introduce errors to the initial program extraction process. This also reduces the total number of scenes. in our data split.
Fastsynth (Ritchie et al., 2018) was not originally trained on 3D-FRONT. It was also trained with object ordering. We retrained Fastsynth on the same data splits of 3D-FRONT as our method, and also without object ordering for a fair comparison with ATISS.
Appendix G More mask examples
We show more examples of masks produced by our and baseline methods. These are shown in Figures 14, 15, 16, 17, and 18.
Figure 14. Additional Mask Examples
[…]
TODO
Figure 15. Additional Mask Examples
[…]
TODO
Figure 16. Additional Mask Examples
[…]
TODO
Figure 17. Additional Mask Examples
[…]
TODO
Figure 18. Additional Mask Examples
[…]
TODO
Appendix H More scene examples
We show more examples of scenes generated by our method and baseline methods. These are shown in Figures 19 and 20.
Figure 19. More examples of scenes generated by our and baseline methods
[…]
TODO
Figure 20. More examples of scenes generated by our and baseline methods
[…]
TODO

### v2

@@ Appendix B Further Details on Constraints
Figure 12: Constraint Examples: Shown are examples of constraints, their input scene and object, and their executed masks. Objects are colored by how they appear in the scene visualization. The original scene shows where the query object was originally placed. There are four masks for each constraint. Each mask represents a possible orientation of the query object. For example the align constraint contains placement options for only one orientation (the orientation of the reference wall object). The other constraints contain placement options for all possible orientations, but those locations are constrained based on the input arguments.
@@
We assume the input objects to each constraints have the following assumed properties. Objects in each category are aligned in a canonical coordinate frame. They must also come labeled with whether they are meant for holding humans (i.e. beds and chairs). Objects and their original scenes must also come with semantically meaningful sizes, scales, and distances since geometric heuristics described for each constraint are based on physically meaningful quantities such as the average reaching distance. Objects and scenes which satisfy this criteria come from preprocessing the 3DFRONT [8] dataset. Visualizations of each constraint is shown in Figure 12.
Appendix C Initial Program Extraction
As a pre-processing step, scenes with major inter-object bounding box collisions are removed from the dataset. These scenes can produce errors in the scene extraction process.
Appendix D Object Placement Classifier Architecture
We perform experiments on two architectures of object placement classifiers. We test a CNN based classifier that uses a top-down image representation  [30] with an additional input channel denoting the query object in the scene. We also test a transformer-based classifier that uses the object encoder described in Section 5 with an additional learned vector embedding added to the query object vector encoding. While our CNN based classifier reports higher precision (accepts fewer invalid programs), and results in better downstream quantitative metrics, we report results using the transformer classifier due to constraints on computation. The transformer model is computationally cheaper than the CNN because it does not require rasterization of the top-down view of the scene.
Appendix E Scene annotation software
Figure 13: A screen shot of our annotation software. On the left shows a partial scene and the object hovers on the mouse pointer. Users can draw rectangles to visualize possible placements, and then confirm them once confident in their drawing.
We built a browser based scene annotation software to facilitate the annotation of partial scenes, objects, and where they could go. We recruited 15 university students and young working professionals to participate in the annotation. Participants were given partial scene and object and asked to mark all the possible places that object can go. No time limit was enforced, but users spent an average of 18 seconds on each partial scene and object, or 30 minutes in total for 100 scenes.
The scene annotation software, built in react, allows users to draw rectangles denoting possible centroid locations of the object for a given orientation. Users can visualize what these proposed placements look like in the scene before confirming these placements. A screenshot of this software are shown in Figure 13
Appendix F Baselines Details
The retraining of both baseline methods [26, 30] on 3D-FRONT [8] differ slightly from their original training setting. Both baselines are trained on the same scenes our method trains on and both without object ordering.
ATISS’s [26] data preprocessing parses living rooms and dining rooms with a maximum side length of 13.2 meters, and bedrooms and libraries with a maximum side length of 6.2. Our data processing only parses scenes with maximum side of 6.2 for all room types. During our data preprocessing, scenes with significant inter-object penetration are removed from the dataset as they introduce errors to the initial program extraction process. This also reduces the total number of scenes. in our data split.
Fastsynth [30] was not originally trained on 3D-FRONT. It was also trained with object ordering. We retrained Fastsynth on the same data splits of 3D-FRONT as our method, and also without object ordering for a fair comparison with ATISS.
Appendix G Edge Attention mechanism
@@
Figure 14: In this example, we want to place a nightstand in a room that already contains a bed and a nightstand on the left side of it. The program which specifies this placement should always predict placement on the right side of the bed. Without edge attention the program will incorrectly place the nightstand on the left side 50% of the time. With edge attention attention our model correctly attends to the spatial relationships between objects in the scene and predicts placement on the right side every time.
@@
Consider the setting in Figure 14. The partial scene contains a single bed and nightstand to the right of it. The query object is also a nightstand. Given this setting our generative model should predict a program that places the nightstand on the left of the bed 100 % of the time. Instead, the logits corresponding to which side of the bed the program will place the nightstand is split 50-50 left and right. This is likely due to nightstands appearing on either side of the bed with equal frequency. We find this is empirically true on both a model trained on real scene data and a toy setting containing just beds and nightstands.
@@
$Z=\text{Softmax}(\frac{W_{q}X(W_{k}X)^{T}}{\sqrt{d_{k}}})W_{v}X$ (2)
@@
For each object, denoted by its index $i$, we compute key and value vectors $K$’ and $V$’ with respect to it. The attention weights are computed as $QK^{T}+QK’^{T}$, where the $QK’^{T}$ term acts as a correction weight to the original $QK^{T}$ attention weights. The weighted sum of $V$’ using these attention weights is added to the normal output of attention for the object. Given edge values $E_{i}\in\mathbb{R}^{txd}$, our edge attention mechanism adds inter object information to each output of the original attention mechanism $Z_{i}$.
$Z_{i}=Z_{i}+\text{Softmax}(\frac{W_{q}X(W_{k}X)^{T}+W_{q}X(W_{ek}E_{i})^{T}}{\sqrt{d_{k}}})W_{ev}E_{i}$ (3)
Figure 15: More examples of annotated masks alongside the location distributions predicted by our method and other methods.
Figure 16: More examples of annotated masks alongside the location distributions predicted by our method and other methods.
Figure 17: More examples of annotated masks alongside the location distributions predicted by our method and other methods.
Figure 18: More examples of how less training data affects the predicted location distributions of our method and baselines
Figure 19: More examples of scenes generated by our and baseline methods
Figure 20: More examples of scenes generated by our and baseline methods